OpenStudy (anonymous):

Simplify: 3/7 - 1/7 + 5/7 - 9/7

OpenStudy (anonymous):

same denominator=> 3-1+5-9=-2 -2/7

OpenStudy (anonymous):

Show work

OpenStudy (anonymous):

2.285714286 well thats what my calculator told me, i did (3/7-1/7=2/7) (2/7+5/7= 7/7=1 (1-9/7=2.285714286